OpenStudy (anonymous):

2x-8 is congruent to x+17 solve for x

OpenStudy (anonymous):

equate the two and solve for x 2x - 8 = x + 17 Can u manage frm here???

OpenStudy (lilg132):

2x - 8 = x + 17 so we need to get the variables all on one side and the numbers on the other so what we do is 2x - x = 17 + 8 (remember bringing it to the opposite side will change it to the opposite of what it is e.g. -8 to the other side will be +8) so now we simplify this equation and we get this : x = 25 dont forget to give me a medal or a good answer :D

OpenStudy (anonymous):

what is 2x-8 plus x+17